OpenStudy (anonymous):

A building lot in a city is shaped as a 30° -60° -90° triangle. The side opposite the 30° angle measures 41 feet. a. Find the length of the side of the lot opposite the 60° angle. b. Find the length of the hypotenuse of the triangular lot. c. Find the sine, cosine, and tangent of the 30° angle in the lot. Write your answers as decimals rounded to four decimal places.

OpenStudy (nirmalnema):



OpenStudy (nirmalnema):

let the side opposite to 60 deg. be x tan60 = perpendicular/base \[\sqrt{3}=x/41 \]

OpenStudy (nirmalnema):

\[x=41\sqrt{3}\] =1.73*41=71.01

OpenStudy (nirmalnema):

2) hypo =\[\sqrt{per^2+base^2}\]

OpenStudy (nirmalnema):

=81.9 approx 90

OpenStudy (nirmalnema):

3) sine30=1/2=0.5 cosine30 = rt 3 /2=0.8 tan30 =1/rt3

OpenStudy (nirmalnema):

@amber.smith19 do u understand